OpenStudy (anonymous):

For the acute angle 0(theda) with cot 0= 1 , find the following. Give exact answers; NO DECIMALS. If the answer involves a square root it should be entered as sqrt . E.g. the square root of 2 should be written as sqrt(2)). sin0= cos0= tan0= sec0= csc0=

OpenStudy (anonymous):

tan = 1 sin = sqrt2/2 cos = sqrt2/2 sec = sqrt2 csc = sqrt2 all positive if in quadrant one